what is the ordered pairs
x-2y=-2
x+2y=10

Found 2 solutions by josgarithmetic, TeachMath:
Answer by josgarithmetic(39620)

Which kind of "the ordered pair"?

2y-2=10-2y
4y-2=10
4y=12
y=3
Can you decide what to do next?

Answer by TeachMath(96)

Add both to get: 2x = 8
x = 4
substitute 4 for x to get: 4 + 2y = 10
2y = 6
y = 3